OpenStudy (anonymous):

solve the equation using factoring or by finding square roots x^2-49=0

OpenStudy (anonymous):

The answer is (x+7)(x-7)

OpenStudy (anonymous):

x^2=49 x=+7,-7 (two rots)

OpenStudy (anonymous):

x^2 - 49 = 0 x^2 = 49 Take square root of both sides. sqrt{x^2} = sqrt{49} Answer: x = 7 and x = -7 When taking the square root, you will always get a positive and negative answer.

OpenStudy (anonymous):

thank you